1. Evaluate the mean and standard deviation for the following.

(a) The young modulus^(GN/m^2) for stainless steel is given by:


Young modulus 193 194 195 196 197 198 199 200

Frequency 7 7 10 19 13 9 8 3


B) The tensile strength of the aluminum alloy is given by:

Tensile strength T / frequency

320^ ≤ T < 350 4

350^ ≤ T < 380 12

380^ ≤ T < 410 18

410^ ≤ T < 440 16

440^ ≤ T < 470 15

470^ ≤ T < 500 10

500^ ≤ T < 530 3

530^ ≤ T < 560 7


1
Expert's answer
2021-04-11T16:52:40-0400

a)



Sample size "N=76."

"\overline{x}=\frac{1}{N}\sum x_in_i\\
x_i\text{ --- value (Young's modulus)}\\
n_i\text{ --- corresponding frequency}\\
\overline{s}=\sqrt{\frac{1}{N}\sum (x_i-\overline{x})^2n_i}"

b)



Sample size "N=85."

"\overline{p_n}(x)=
\begin{cases}
\frac{l_j}{N\cdot h}, x\in [a+jh,a+(j+1)h)\\
0,\ x\notin (a,b)
\end{cases}\\
l_j\text{ --- number of values in the interval } [a+jh,a+(j+1)h)"

"h=30"

"x_i^*=\frac{x_i+x_{i-1}}{2}\\
\overline{s}=\sqrt{\frac{1}{N}\sum(x_i^*)^2n_i-(\frac{1}{N}\sum x_i^*n_i)^2}\\
\overline{x}=\frac{1}{N}\sum x_i^*n_i"